OpenStudy (anonymous):

lim x->1 (x^25-1)/(x-1), determine if the limit exists and if it does find its value.

OpenStudy (zzr0ck3r):

have you learned lhopita rule yet?

OpenStudy (zzr0ck3r):

l'hopital*

OpenStudy (anonymous):

i have no idea what that is

OpenStudy (zzr0ck3r):

its x^(25)?

OpenStudy (anonymous):

yeah in the problem its x^25

OpenStudy (anonymous):

\[\frac{x^{25}-1}{x-1}=1+x+x^2+x^3+x^4+x^5+x^6+x^7+x^8+x^9+x^{10}+x^{11}+x^{12}+x^{13}+ \]\[x^{14}+x^{15}+x^{16}+x^{17}+x^{18}+x^{19}+x^{20}+x^{21}+x^{22}+x^{23}+x^{24} \]There are 25 terms. 25*1 = 25

OpenStudy (zzr0ck3r):

wooo sweet, can you explain that a little more?

OpenStudy (zzr0ck3r):

Im thinking if he has not learned l'hopital rule yet, then he shuold be able to do it another way.

OpenStudy (anonymous):

The RHS came from Mathematica's "Apart" function. 1^n where n is any positive integer is equal to 1

OpenStudy (zzr0ck3r):

very cool. what are you doing in class @kjblasion ? are you finding limits by approximation methods, graphically, or are you finding them with algebra only?

OpenStudy (anonymous):

oh that makes sense.. but in my class they are making us find the limits through the derivative.

OpenStudy (zzr0ck3r):

if you take the derivative of the top and bottom as many times as it takes to get it to a form that is determenant, then plug in 1

OpenStudy (zzr0ck3r):

this is called l'hopital rule.

OpenStudy (zzr0ck3r):

so you have (25x^24)/1 when x =1 you get 25

OpenStudy (anonymous):

L'Hôpital's rule http://en.wikipedia.org/wiki/L'H%C3%B4pital's_rule Divide the derivative of the Numerator by the derivative of the Denominator.\[\frac{\left(25 x^{24}\right)}{1} \] The limit as x->1 with this method is also 25.

OpenStudy (anonymous):

I have seen at least one limit problem presented on this site where L'Hôpital's rule was not allowed to be used in the solution.

OpenStudy (anonymous):

Oh wow, that makes more sense i was trying to take the quotient rule and I'm pretty sure i screwed up some where. Thanks for helping me solve the problem!!!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

differentiate both the denominator and numerator w.r.t. x until u get the expression free from the form of 0/0 .... then u can put the limits... which will give the same result.....